OpenStudy (anonymous):

The perimeter of a photograph is 32 inches. It is 6 inches wide. How long is it?

OpenStudy (anonymous):

no one knows the answer?:(

OpenStudy (anonymous):

ok time out

OpenStudy (anonymous):

10 inches

OpenStudy (anonymous):

10, because 32-12 = 20 then divide it by 2 so it is 10

OpenStudy (anonymous):

since a photograph is square or rectangular, it has 2 parallel pairs of sides which are both equal in length. So it's width being 6 inches, means that there are 2 sides that are 6 inches long. So we have 32 - 2(6) = 2x 32 - 12 = 2x 20 = 2x 10 = x SO it is 10 inches long.